OpenStudy (anonymous):

find x intercepts of y=2(x-1)^2-8. I keep getting +/-3 but that's wrong, why?

OpenStudy (amistre64):

4*2 - 8 x-1 = 4 x-1 = -4

OpenStudy (amistre64):

ugh ... typoed lol

OpenStudy (amistre64):

x-1 = 2 or -2

OpenStudy (zarkon):

I get -1 and 3

OpenStudy (amistre64):

-3-1 = -4 -4^2 = 16 16-8 not= 0 3-1 = 2

OpenStudy (anonymous):

my answer says x int. are (3, 0) and (-1, 0). I'm adding the 8, dividing by 2 and square rooting to get x-1=+/-3 but when i plug it in to graph the x intercepts are -1 and 3

OpenStudy (amistre64):

your not doing the last part correctly

OpenStudy (zarkon):

\[2(x-1)^2-8=0\] \[2(x-1)^2=8\] \[(x-1)^2=4\] \[(x-1)=\pm\sqrt{4}=\pm2\] \[x=1\pm2\] thus \[x=-1,3\]

OpenStudy (anonymous):

ahh, got it, so I was just combining them incorrectly

OpenStudy (anonymous):

That's great zarkon,I just plotted the graph and got the same result.

OpenStudy (anonymous):

Thanks for pointing out my mistake zarkom! (I tend to get the overall concept but then make stupid little errors that confuse me.. :) )

OpenStudy (anonymous):

So, to find the x intercept set y = 0 in the equation and then $$ 2 (x - 1)^2 - 8 = 2 (-3+x) (1+x)=0\Rightarrow x=-1,3$$

OpenStudy (zarkon):

\[2(x-1)^2-8\] \[2[(x-1)^2-4]\] \[2[(x-1-2)(x-1+2)]\]...